OpenStudy (anonymous):

f(x) = -4x + 7 and g(x) = 10x - 6, find f(g(x)).

OpenStudy (ipwnbunnies):

f(g(x)) means you plug in g(x) into the x's in f(x)! It's kinda weird!

OpenStudy (anonymous):

I don't understand how to like solve the answer though, can you help me? @iPwnBunnies

OpenStudy (anonymous):

I'm really bad at Algebra. I'm taking Algebra 2 but i'm only n 8th grade

OpenStudy (ipwnbunnies):

Sure. Let me just do the plugging in, I'm sure you can simplify. f(g(x)) = -4(10x-6) + 7. Do you see now? We plugged in the function g(x) into the x's in f(x).

OpenStudy (anonymous):

so then you just solve for x?

OpenStudy (anonymous):

-40x+31?

OpenStudy (ipwnbunnies):

Yes! I don't think you solve for x. You keep it as f(g(x)) = -40x + 31. Good job!

OpenStudy (anonymous):

what about this one? f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (anonymous):

i don't know how to do that first part that you did but after that part i can solve the equation

OpenStudy (ipwnbunnies):

Do the exact same thing I just showed you. I'll be here.

OpenStudy (anonymous):

ok ill try it, when I'm done ill show you

OpenStudy (ipwnbunnies):

Ok/

OpenStudy (anonymous):

so its 9(-x+3)-2

OpenStudy (ipwnbunnies):

Yes! Now expand it, simplify, and set it equal to f(g(x))

OpenStudy (anonymous):

simplified to f(g(x))=-9x+25?

OpenStudy (ipwnbunnies):

Yep. Good work man. :)

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

I've done most of the rest of my problems, would you mind checking them for me? I do online school and i don't really have anyone here to help me

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Ok

OpenStudy (anonymous):

f(x) = x2 - 8x + 5. Find f(-1).

OpenStudy (anonymous):

and i got 14? @iPwnBunnies

OpenStudy (ipwnbunnies):

Correct.

OpenStudy (anonymous):

f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). and i got -3

OpenStudy (ipwnbunnies):

Yes, correct!

OpenStudy (anonymous):

f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). and i got -7

OpenStudy (ipwnbunnies):

Correct.

OpenStudy (anonymous):

and now this one I'm having trouble with as well

OpenStudy (anonymous):

f(x) = 2x - 6 Solve f^-1(x) when x = 2.

OpenStudy (ipwnbunnies):

f^(-1)(x) is the inverse of f(x). What do you is change places of y and x in f(x). Solve for y again.

OpenStudy (ipwnbunnies):

What you do*

OpenStudy (anonymous):

wait I don't understand, can you give me an example? I'm sorry

OpenStudy (ipwnbunnies):

Sure. For example, we have f(x) = 3x + 2. To find the inverse, rewrite f(x): y = 3x + 2 2. Change the places of y and x: x = 3y + 2 3. Solve for y, and replace it with f^(-1)(x): y = (x-2)/3 = f^(-1)(x)

OpenStudy (anonymous):

ok… so the answer to the one i was doing, is 4 correct?

OpenStudy (ipwnbunnies):

Correct, great job man.

OpenStudy (anonymous):

ok, thank you! Last one i promise! haha f(x) = 2x + 2. Solve f-1(x) when x = 4

OpenStudy (ipwnbunnies):

Do the same process I showed you to get the inverse of f. Plug in 4 into the inverse.

OpenStudy (anonymous):

so the answer would be one?

OpenStudy (ipwnbunnies):

Yup.

OpenStudy (anonymous):

thank you so much!!! your the best!!!

OpenStudy (ipwnbunnies):

No prob.